# Smooth-edged Perturbations Improve Perturbation-based Image Explanations

### v1

@@ IV Discussion
This work shows that the smooth-edged masks used in the original RISE implementation can be modified to work with many different attribution pipelines and that this improves performance on occlusion metrics. However, occlusion metrics do not necessarily correlate with usefulness to humans. It may be the case that per-pixel attribution simply gives advantages in performance calculation that are not noticeable in user testing, for which further work is needed.
@@
The results also show that each part of the pipeline that is explored can have a significant impact on performance. Most works that introduce some form of perturbation-based image explanations often introduce an entire pipeline but do not examine the parameters of that pipeline separately. This leads to a poor understanding of what makes one method better, especially when later works compare those pipelines against each other [25, 26]. Contrastingly, this work along with Blücher et al., [8] shows how the different parameters can be analyzed independently.
The evaluation in this work relied on the explanation methods being separable into different parameters that could be combined in various ways. This is not always the case, even if the method otherwise produces sound explanations. For example, the original RISE shifts the perturbation masks by some pixels so as not to center the same pixels every time. This approach works with the RISE attribution method since it can directly assign influence to pixels. However, this is not feasible for other methods, as such shifting could not be evaluated with the experiments conducted in this work. Additionally, the use of occlusion metrics requires attribution scores for individual features. For example, the CIU method can be used when combinations of features are perturbed simultaneously, however, those explanations instead give attribution to how beneficial the combinations are, rather than splitting the influence between the features. Another example is using decision trees as surrogate models. Decision trees are typically interpretable but do not assign influence to features directly.
@@
A general issue with all current perturbation-based methods is that they require that the model be run multiple times. This inevitably scales the computation needed by at least a factor equal to the sample size used. With ever-increasing computational demands by newer DNN models, even a low sample size let alone thousands of samples, may be unrealistic to presume for an explanation of a single decision. Developing perturbation-based methods that can give good explanations with low sample size is therefore a promising future direction. In some cases, such as medical diagnosis prediction, the need for and the value of explanations are likely high enough that it is worth increasing computational demands by factors of thousands.
One contender to perturbation-based methods is gradient-based methods. Gradient-based post-hoc methods utilize that DNNs are typically differentiable and use the gradients of the prediction to calculate an explanation. This gives gradient-based methods the advantage that they often do not need multiple calls to the model. However, gradient-based explanation methods, especially in the computer vision domain, have multiple times been shown to be unreliable [27, 28, 29]. Perhaps a combination of the different post-hoc paradigms could benefit from the reliability of perturbation-based methods and the lower computational demands of gradient-based methods. For example, the initial gradient-based explanation could inform the optimal segments or samples to use with a permutation-based approach.
Ultimately, the true measure of any explanation is its usefulness to humans. For example, a prior study found that users preferred CIU explanations to LIME and SHAP [20], which is not obvious from the results in this work. However, the number of different parameter combinations that exist in XAI is too many for human evaluators. As such, future works might strive to use metrics such as SRG to find the best candidate pipelines and then compare those using human evaluation. Such experiments would require an additional step to the pipeline; communicating. How an explanation is communicated to humans can vary between implementations and is another factor that can disrupt experiments. As such a study focusing solely on communication and visualization of image attribution would be beneficial.

### v2

@@ IV Discussion
Most works that introduce or compare perturbation-based image explanation methods do not examine the parameters of the explanation pipelines separately. This leads to a poor understanding of why a method performs well and can lead to questionable claims of some methods being better than others [28, 29]. The results of this work show that each parameter of the perturbation-based image explanations pipeline significantly impacts performance.
From the results, it is clear that using SLIC as opposed to grid segmenting improves performance significantly. Per-pixel attribution, which requires smooth masks, also improves performance compared to per-segment attribution. Finally, using SLIC with per-pixel attribution further improved performance. This shows that using Gaussian filters or other alternatives to bilinear smoothing is promising as it allows the use of better-performing segmentation methods. The choice of attribution calculation has a noticeably smaller impact on performance when compared to the choice of segmentation method or the difference between per-segment and per-pixel attribution. It is interesting that PDA and RISE attribution perform differently when they perform equally in the ideal case. It seems that, with a limited sample size, using the average decrease in prediction when a segment is excluded (PDA) is less stable than using the average increase when it is (RISE).
@@
It is clear that an increased sample size does improve performance. However, the computation needed scales directly with the sample size. With ever-increasing computational demands by newer DNN models, even a low sample size let alone thousands of samples, may be unrealistic to presume for an explanation of a single decision. Developing perturbation-based methods that can give good explanations with low sample size is therefore a promising future direction. In some cases, such as medical diagnosis prediction, the need for and the value of explanations are likely high enough that it is worth increasing computational demands by factors of thousands.
The evaluation in this work relied on the explanation methods being separable into different parameters that could be combined in various ways. This is not always the case, even if the method otherwise produces sound explanations. For example, the original RISE implementation shifts the perturbation masks by some pixels so as not to center the same pixels every time. This approach works with the RISE attribution calculation since it can directly assign influence to pixels. However, this is not feasible for other methods, as such shifting has not been used in this evaluation. Additionally, the use of SRG and similar metrics requires attribution scores for individual features. For example, the CIU method can use intermediate concepts to explain combinations of features, which results in explanations for the combinations rather than the individual features.
This work shows how different parameters affect performance as measured by SRG. SRG measures how well the explanation reflects the actual impact of the attributed regions on the model. While ensuring that attribution maps explain the model correctly is helpful, it is not known whether it correlates with usefulness to humans. Existing evaluation on whether proxy metrics correspond to usefulness for humans is limited and requires further investigation [13]. It is also questionable whether proxy metrics can be used to compare different types of attribution maps at all. Attribution maps that indicate large regions, that provide heatmaps, or that show important edges and textures are likely to be interpreted differently by humans in a way not reflected by most proxy metrics.